OpenStudy (anonymous):

given x^2+b=0, state a value of b that would make A)x a real number B)x an imaginary number C)x a complex number

OpenStudy (anonymous):

what is that

OpenStudy (daniellelovee):

-1 is an imaginary number since it equals i^2 @shamim you were on the right path

OpenStudy (shamim):

x^2+b=0 x^2=-b x=√(-b) x=√-1*√b x=i*√b

OpenStudy (shamim):

Is it readable

OpenStudy (shamim):

Anyway i m frm android phone n watching some ? Sign

OpenStudy (daniellelovee):

so this would be imaginary right?

OpenStudy (daniellelovee):

@shamim?

OpenStudy (daniellelovee):

@helpme375 what did you get?

OpenStudy (shamim):

Ya u r right

OpenStudy (shamim):

Complex number a real part n a imaginary part